OpenStudy (anonymous):

PLEASE HELLLLLLLLLLLLLLLLLP |x+4|^2-10|x+4|=24

OpenStudy (anonymous):

need help exspanding love

OpenStudy (anonymous):

\[|x+4|^{2}-10|x+4|=24\]

OpenStudy (anonymous):

|x+4|^2 = (x+4)^2 = x^2 + 8x + 16

OpenStudy (anonymous):

are you sure you distribute?

OpenStudy (anonymous):

because \(\left(x+4\right)^2 \ge 0\)

OpenStudy (anonymous):

hes is right

OpenStudy (mathmath333):

u can also put |x+4| =a and then solve the quadratic

OpenStudy (anonymous):

so\[|x+4| = \frac{ \left( x+4 \right)^2-24 }{ 10 }\] you could graph or solve for the 2 cases... \(x \ge -4\) and \(x < -4\)

OpenStudy (anonymous):

x will have to be > -1 for first case and x < -5 for second case.